Assume that the population of heights of male college students is approximately normally distributed with mean m of 70 inches and standard deviation s of 4.15 inches. Show all work.
(A) Find the proportion of male college students whose height is greater than 68 inches.
(B) Find the proportion of male college students whose height is no more than 68 inches.
I have my work, but I want to make sure I am doing it right. We are using z score tables to determine the area(probability) based on the information given. The only thing I am really uncertain of is the on B where it is not just less than, but also = to 68 in.

Answer by stanbon(75887)

Assume that the population of heights of male college students is approximately normally distributed with mean m of 70 inches and standard deviation s of 4.15 inches. Show all work.
(A) Find the proportion of male college students whose height is greater than 68 inches.
z(68) = (68-70)/4.15 = -0.4819
P(x > 68) = P(z > -0.4819) = 0.6851
-------------------------------------
(B) Find the proportion of male college students whose height is no more than 68 inches.
Ams: 1 - 0.6851 = 0.3159
=================================
I have my work, but I want to make sure I am doing it right. We are using z score tables to determine the area(probability) based on the information given. The only thing I am really uncertain of is the on B where it is not just less than, but also = to 68 in.
Comment: The end values make no difference.
Example: P(x > 68) is the same as P(x >= 68)
=================================================
Cheers,
Stan H.
=====================